OpenStudy (anonymous):

What is the radian measure of an angle that is 105 degrees 32' 45'' ?

OpenStudy (ranga):

convert 105 degrees 32' 45'' to degrees first. 3600'' = 1 degree. 45'' = 45/3600 degree 60' = 1 degree 32' = 32/60 degree 105 degrees 32' 45'' = 105 + 32/60 + 45/3600 degrees. Simplify. (1) 180 degrees = pi radians x degrees = x * pi/180 radians. Multiply what you get in step (1) by pi/180 to get convert to radians.

OpenStudy (anonymous):

so its false?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

I don't think this is a true or false question. It is asking you to convert from degrees to radians. 105 degrees 32' 45'' = 105 + 32/60 + 45/3600 = 105.5458 degrees To convert degrees to radians, multiply by pi/180: 105.5458 * pi / 180 = 1.8421 radians.